# Galaxy Spin Alignment with Tidal Fields in the SDSS-IV MaNGA Survey

### v1

@@ Abstract
The tidal torque theory (TTT) predicts that galaxy spins are correlated with the surrounding tidal field, reflecting how angular momentum is acquired during structure formation. We present a new observational test of this prediction using the final data release of the SDSS-IV MaNGA integral field spectroscopy survey, which enables direct spin measurements from stellar and ionized gas kinematics for nearly 10,000 galaxies. Unlike previous studies that relied on filament orientations, we directly utilize the three-dimensional tidal field reconstructed from the galaxy distribution, providing a physically defined reference frame for the analysis. We find that massive galaxies tend to align their spins parallel to the intermediate axis of the tidal field, consistent with the prediction of the TTT. They also tend to orient perpendicular to the major axis, hinting at the possible influence of baryonic processes. In contrast, low-mass galaxies exhibit the opposite trend, with a transition mass of $M_{*}\sim 10^{10}$$10^{10.5}\,M_{\odot}$. No significant alignment is detected with respect to the minor axis across all stellar masses. We further examine the dependence on morphology and environment, finding that S0 and early-type spiral galaxies exhibit stronger alignment signals than late-type spirals. The alignment trend becomes particularly pronounced in regions of high tidal anisotropy and high overdensity. A mutual information analysis identifies these environmental factors as the dominant drivers of the observed trends. Our results provide the most direct empirical evidence to date for the connection between galaxy spins and the cosmic tidal field.
@@
journal: ApJLfacilities: Sloan(MaNGA)software: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Marvin (B. Cherinka et al., 2019), Matplotlib (J. D. Hunter, 2007), NumPy (C. R. Harris et al., 2020), PaFit (D. Krajnović et al., 2006)
show]jsmoon.astro@gmail.com
@@ 1 Introduction
Galaxy spins are not randomly oriented but exhibit preferred orientations in space. The origin of the galaxy spin alignment can be understood within the framework of the tidal torque theory (TTT), which suggests that a misalignment between a protohalo’s inertia tensor and the surrounding tidal field generates a torque during the early stages of cosmic evolution (P. J. E. Peebles, 1969; A. G. Doroshkevich, 1970; S. D. M. White, 1984). In the linear TTT, galaxy spins are expected to align preferentially with the intermediate axis of the local tidal tensor (J. Lee & U.-L. Pen, 2000; B. M. Schäfer, 2009). As structure formation proceeds, the direction of these spins can be altered by nonlinear vorticity in the cosmic flow (e.g., N. I. Libeskind et al., 2013; C. Laigle et al., 2015) and by mergers and tidal encounters between galaxies (e.g., P. R. Capelo & M. Dotti, 2017; J.-S. Moon et al., 2021). Nevertheless, the spin directions remain intimately connected to the geometry of the cosmic web, tracing how angular momentum builds up within the evolving large-scale structure (e.g., S. Codis et al., 2012). Understanding how galaxy spins are coupled to their large-scale environments is therefore crucial for disentangling the competing physical processes that drive the acquisition and evolution of angular momentum in galaxies.
@@
Confirming the spin alignment observationally has long been challenging due to limited data availability. Early attempts inferred spin orientations indirectly from galaxy shapes (J. Lee & P. Erdogdu, 2007; Y. Zhang et al., 2015; I. Pahwa et al., 2016). However, the alignment between the shape and rotation axes holds only for galaxies that are predominantly rotation-supported. Defining the cosmic web adds further complications. Reconstructing the tidal field from galaxy distributions is non-trivial, and most observational studies have instead relied on filament orientations (B. J. T. Jones et al., 2010; E. Tempel et al., 2013; E. Tempel & N. I. Libeskind, 2013; A. Hirv et al., 2017), identified as density ridges in the galaxy distribution (e.g., T. Sousbie, 2011; Y.-C. Chen et al., 2015; E. Tempel et al., 2016). These approaches are sensitive to the details of filament detection and to redshift-space distortions—particularly the finger-of-god effect—which introduces significant systematic uncertainties (see, e.g., E. Tempel et al., 2013; C. Welker et al., 2020).
@@
The advent of integral field spectroscopy (IFS) has transformed this field by allowing direct spin measurements from spatially resolved stellar kinematics. Using the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA) survey (K. Bundy et al., 2015), A. Krolewski et al. (2019) conducted the first measurement of spin alignments with nearby filaments but found no significant signal. Building on this work, K. Kraljic et al. (2021) analyzed 4,633 MaNGA galaxies from the Sloan Digital Sky Survey (SDSS) Data Release 15 (D. S. Aguado et al., 2019) and measured three-dimensional spin–filament alignments, finding parallel alignments for spirals and perpendicular ones for S0s. Using the Sydney–AAO Multi-object Integral-field spectrograph (SAMI) Galaxy Survey (S. M. Croom et al., 2021), C. Welker et al. (2020) detected a mass-dependent spin transition with respect to the filament spine. S. Barsanti et al. (2022) demonstrated spin alignment signals in about 3,000 SAMI galaxies, showing that the transition correlates most strongly with bulge mass. S. Barsanti et al. (2023) subsequently investigated how black hole activity relates to the spin alignment.
@@
In this Letter, we present a new test of galaxy spin alignment with respect to the local tidal field using the latest IFS data from the MaNGA survey. This is the first study to examine the spin alignment with respect to the tidal tensor using direct spin measurements derived from spatially resolved kinematics. This approach offers a distinct advantage for a physically well-motivated and consistently defined measure of the local gravitational deformation, applicable even in regions far from filament spines. Our analysis uses the final MaNGA data release (Abdurro’uf et al., 2022), which contains nearly 10,000 galaxies, nearly doubling the sample size of previous IFS-based studies and enabling a statistically robust detection of alignment signals. These advances allow us to probe the intrinsic connection between galaxy spin and the underlying cosmic tidal field, offering a new window into the physics of galaxy formation within the cosmic web.
@@ 2.1 MaNGA Galaxies
We utilize the IFS data from the complete release of the MaNGA survey, which is part of SDSS Data Release 17 (K. Bundy et al., 2015; Abdurro’uf et al., 2022). All MaNGA data can be accessed via a Python-based tool called Marvin (B. Cherinka et al., 2019). MaNGA employs a set of integral field units, each consisting of between 19 and 127 fibers, to obtain spatially resolved spectra of nearby galaxies with a spectral resolution of $R\sim 2000$ and a wavelength coverage of 3600–10400 Å. The raw spectra are flux-calibrated and combined into three-dimensional data cubes through the MaNGA Data Reduction Pipeline (DRP; D. R. Law et al. 2016, 2021), and these data cubes are then processed by the MaNGA Data Analysis Pipeline (DAP; K. B. Westfall et al. 2019; F. Belfiore et al. 2019) to derive various astrophysical quantities, including stellar and gas kinematics, emission-line fluxes, and spectral indices. The DAP performs full-spectrum fitting on the binned spectra, providing the line-of-sight velocity fields for both the stellar and ionized gas components.
@@
Our sample includes all galaxies observed in MaNGA, except those from four ancillary programs targeting non-galactic objects (DEEP_COMA, IC342, M31, and GLOBULAR_CLUSTER). We also exclude data flagged by the DAPQUAL bitmask as critical failures in either the DRP or DAP. Our parent sample consists of 10,058 unique galaxies, after removing duplicate entries in MaNGA. In this study, we include only disk galaxies (S0 and spiral types) classified in the MaNGA Visual Morphology Catalogue (MVM-VAC; J. A. Vázquez-Mata et al. 2022), resulting in a sample of 8,199 galaxies.
@@
We measure the position angles of the kinematic major axis, $\phi_{\mathrm{kin}}$, of MaNGA galaxies using the PaFit package (D. Krajnović et al., 2006), which determines the axis that maximizes the match between the velocity field and its bi-antisymmetry model. The central positions and redshifts of the galaxies are adopted from the MaNGA catalog, which for most galaxies is identical to the NASA–Sloan Atlas (NSA; M. R. Blanton et al. 2011). The MaNGA target selection strategy assigns 63 % of the main sample to reach 1.5 effective radii ($R_{e}$) and 37 % to reach $2.5\,R_{e}$ (K. Bundy et al., 2015). For our analysis, we consistently use an elliptical aperture with a semi-major axis of $1.5\,R_{e}$ for each galaxy, with the aperture parameters defined by the NSA elliptical Petrosian photometry. We also exclude all bad pixels with inverse variance set to zero from the measurement.
@@
The spin axis is defined as perpendicular to the kinematic major axis. The vector direction is determined by identifying the approaching and receding sides from the inverse-variance-weighted averages of line-of-sight velocities on either side of the spin axis. Following the method of J. Lee & P. Erdogdu (2007) and J. Lee (2011), we compute the three-dimensional (3D) unit spin vector, $\hat{\boldsymbol{j}}$, in spherical coordinates as
@@
where $\vartheta$ denotes the position angle of the projected spin vector, measured eastward from north, ranging from 0 to $2\pi$. The inclination angle $i$ is estimated under a circular disk approximation from the observed axial ratio $(b/a)$ and the intrinsic axial ratio $p$, and is given by
@@
where we simply adopt $p=0.158$ to be consistent with K. Kraljic et al. (2021).
@@
with $\alpha$ and $\delta$ representing the right ascension and declination of each galaxy, respectively. Due to the sign ambiguity of the radial component $\hat{j}_{r}$, the 3D spin vector cannot be uniquely determined without additional assumptions, resulting in a two-fold degeneracy (J. Lee, 2011; P. Motloch et al., 2021). While many previous studies arbitrarily selected one of the two vectors (e.g., K. Kraljic et al., 2021; S. Barsanti et al., 2022), we account for both in our analysis, which differ only by the sign of $\hat{j}_{r}$. We note that our results are not sensitive to either the sign of $\hat{j}_{r}$ or the value of $p$.
@@ 2.2 Reconstructed Tidal Field
We obtain the tidal field around the MaNGA galaxies from the Galaxy Environment for MaNGA Value Added Catalog (GEMA-VAC; M. Argudo-Fernández et al. 2015; Abdurro’uf et al. 2022). The GEMA-VAC provides a characterization of large-scale structures based on the Exploring the Local Universe with the reConstructed Initial Density field (ELUCID) project (H. Wang et al., 2016). The present-day density field within the SDSS volume is reconstructed from the distribution of galaxy groups using the halo-domain method (H. Wang et al., 2009, 2013). The survey volume is partitioned into halo domains, and the mass distribution in each domain is modeled with $N$-body simulations. Redshift-space distortions are iteratively corrected using the line-of-sight component of the reconstructed linear velocity field.
@@
The density field is smoothed with a Gaussian kernel with a comoving scale of 2 $h^{-1}$Mpc. The three eigenvalues of $T_{ij}$, $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$, are obtained by diagonalizing $T_{ij}$ at the position of each galaxy. The corresponding eigenvectors are denoted by the major ($\hat{\boldsymbol{e}}_{1}$), intermediate ($\hat{\boldsymbol{e}}_{2}$), and minor ($\hat{\boldsymbol{e}}_{3}$) principal axes of $T_{ij}$, respectively. The major principal axis, $\hat{\boldsymbol{e}}_{1}$, corresponds to the direction of maximum matter compression. The eigenvalues can be used to classify the large-scale environment based on the number of positive eigenvalues: three positive eigenvalues indicate a cluster environment, two indicate a filament, one indicates a sheet, and none indicates a void. In our sample, a total of 6,376 galaxies have corresponding tidal field data available for analysis.
@@ 2.3 Alignment Angles
The spin–tidal alignment is quantified as the cosine of the angle, $|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|$ ($k=1,2,3$), where $\hat{\boldsymbol{j}}$ and $\hat{\boldsymbol{e}}_{k}$ represent the unit spin vector of each galaxy and the three principal axes of the local tidal tensor, respectively. In three dimensions, a random orientation yields an average of $|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|=0.5$, while parallel and perpendicular orientations correspond to $|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|=1$ and 0, respectively. Since the line-of-sight component of spin vectors cannot be uniquely determined in this study, each galaxy has two possible 3D spin orientations. In the edge-on view ($i=\pi/2$), the two vectors appear identical, while in the face-on view ($i=0$), they appear to point in opposite directions. To account for this ambiguity, we take the average of $|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|$ from the two possible spin vectors to obtain a single alignment metric for each galaxy. The detection significance of the observed alignment is estimated from the 1,000 random permutation samples in which the sky positions ($\alpha$ and $\delta$) of galaxies are randomly shuffled within the sample. For each random permutation, the 3D spin vectors are recalculated using the new sky coordinates through Equations (1) to (7). This approach retains as much orientation information as possible while ensuring that the two spin vectors from the same galaxy are not treated as independent measurements.

### v2

@@ Abstract
The tidal torque theory (TTT) predicts that galaxy spins are correlated with the surrounding tidal field, reflecting how angular momentum is acquired during structure formation. We present a new observational test of this prediction using the final data release of the Sloan Digital Sky Survey IV Mapping Nearby Galaxies at Apache Point Observatory integral field spectroscopy survey, which enables direct spin measurements from stellar and ionized gas kinematics for a sample of 6325 disk galaxies. We utilize the three-dimensional tidal field reconstructed from the galaxy distribution, providing a physically defined reference frame for the analysis. We find that massive galaxies tend to align their spins parallel to the intermediate axis of the tidal field, consistent with the prediction of the TTT, while also showing a tendency to align perpendicular to the major axis. In contrast, low-mass galaxies exhibit the opposite trend, with a transition mass of $M_{*}\sim 10^{10}$$10^{10.5}\,M_{\odot}$. No significant alignment is detected with respect to the minor axis across all stellar masses. We further examine the dependence on morphology and environment, finding that S0 and early-type spiral galaxies exhibit stronger alignment signals than late-type spirals. The alignment trend becomes particularly pronounced in regions of high tidal anisotropy and high overdensity. A mutual information analysis identifies these environmental factors as the dominant drivers of the observed trends. Our results provide new empirical evidence for the connection between galaxy spins and the cosmic tidal field.
@@
facilities: Sloan(MaNGA)software: Astropy (2013A&A...558A..33A; 2018AJ....156..123A; 2022ApJ...935..167A), Marvin (2019AJ....158...74C), Matplotlib (2007CSE.....9...90H), NumPy (2020Natur.585..357H), PaFit (2006MNRAS.366..787K)
@@ 1 Introduction
Galaxy spins are not randomly oriented but exhibit preferred orientations in space. The origin of the galaxy spin alignment can be understood within the framework of the tidal torque theory (TTT), which suggests that a misalignment between a protohalo’s inertia tensor and the surrounding tidal field generates a torque during the early stages of cosmic evolution (1969ApJ...155..393P; 1970Ap......6..320D; 1984ApJ...286...38W; see also 2002MNRAS.332..325P; 2002MNRAS.332..339P; 2020OJAp....3E...3N; 2021MNRAS.502.5528L; 2025MNRAS.543.2222L). In the linear TTT, galaxy spins are expected to align preferentially with the intermediate axis of the local tidal tensor (2000ApJ...532L...5L; 2009IJMPD..18..173S). As structure formation proceeds, the direction of these spins can be altered by nonlinear vorticity in the cosmic flow (e.g., 2013ApJ...766L..15L; 2015MNRAS.446.2744L) and by mergers and tidal encounters between galaxies (e.g., 2007MNRAS.375..489H; 2009MNRAS.398.1742H; 2017MNRAS.465.2643C; 2021ApJ...909...34M). Nevertheless, the spin directions remain intimately connected to the geometry of the cosmic web (e.g., 1996Natur.380..603B; 2008LNP...740..335V; 2024arXiv240716489K), tracing how angular momentum builds up within the evolving large-scale structure (e.g., 2012MNRAS.427.3320C; 2015MNRAS.452.3369C). Understanding how galaxy spins are coupled to their large-scale environments is therefore crucial for disentangling the competing physical processes that drive the acquisition and evolution of angular momentum in galaxies.
Cosmological $N$-body simulations show a mass-dependent trend in spin alignments, with a characteristic transition in the preferred spin orientation around a halo mass of $10^{12-13}\,M_{\odot}$ (2007ApJ...655L...5A; 2007MNRAS.381...41H; 2009ApJ...706..747Z; 2012MNRAS.427.3320C; 2013MNRAS.428.2489L; 2013ApJ...762...72T; 2014MNRAS.440L..46A; 2014MNRAS.443.1090F; 2017MNRAS.468L.123W; 2018MNRAS.473.1562W; 2018MNRAS.481..414G; 2021MNRAS.503.2280G; 2020ApJ...902...22L; 2021MNRAS.502.5528L). However, it remains uncertain whether a similar trend persists for observable galaxies, which probe only the inner regions of their host dark matter (DM) halos. Recent hydrodynamical simulations have predicted that stellar spins also exhibit clear alignment signals, albeit with a somewhat weaker mass-dependent transition than their DM counterparts (2018MNRAS.481.4753C; 2018ApJ...866..138W; 2019MNRAS.487.1607G; 2020MNRAS.493..362K; 2021ApJ...922....6L; 2023ApJ...945...13M).
@@
Confirming the spin alignment observationally has long been challenging due to limited data availability. Early attempts inferred spin orientations indirectly from galaxy shapes (2007ApJ...671.1248L; 2015ApJ...798...17Z; 2016MNRAS.457..695P). However, the alignment between the shape and rotation axes holds only for galaxies that are predominantly rotation-supported. Defining the cosmic web adds further complications. Reconstructing the tidal field from galaxy distributions is nontrivial, and most observational studies have instead relied on filament orientations (2010MNRAS.408..897J; 2013MNRAS.428.1827T; 2013ApJ...775L..42T; 2017A&A...599A..31H), identified as density ridges in the galaxy distribution (e.g., 2011MNRAS.414..350S; 2015MNRAS.454.1140C; 2016A&C....16...17T). These approaches are sensitive to the details of filament detection and to redshift-space distortions—particularly the finger-of-god effect—which introduces significant systematic uncertainties (see, e.g., 2013MNRAS.428.1827T; 2020MNRAS.491.2864W).
@@
The advent of integral-field spectroscopy (IFS) has transformed this field by allowing direct spin measurements from spatially resolved stellar kinematics. Using the Mapping Nearby Galaxies at Apache Point Observatory (MaNGA) survey (2015ApJ...798....7B), 2019ApJ...876...52K conducted the first measurement of spin alignments with nearby filaments but found no significant signal. Building on this work, 2021MNRAS.504.4626K analyzed 4633 MaNGA galaxies from the Sloan Digital Sky Survey (SDSS) Data Release 15 (2019ApJS..240...23A) and measured three-dimensional (3D) spin–filament alignments, finding parallel alignments for spirals and perpendicular ones for S0s. 2025ApJ...987L..30W detected an antiparallel alignment signal between galaxy spins and the spins of their host filaments in MaNGA. Using the Sydney–AAO Multi-object Integral-field spectrograph (SAMI) Galaxy Survey (2021MNRAS.505..991C), 2020MNRAS.491.2864W detected a mass-dependent spin transition with respect to the filament spine. 2022MNRAS.516.3569B demonstrated spin alignment signals in about 3,000 SAMI galaxies, showing that the transition correlates most strongly with bulge mass. 2023MNRAS.526.1613B subsequently investigated how black hole activity relates to the spin alignment.
@@
In this Letter, we present a new test of galaxy spin alignment with respect to the local tidal field using the latest IFS data from the MaNGA survey. This work represents the first study to examine spin alignment by combining (i) a reconstructed tidal field and (ii) kinematically measured spins. The use of the tidal field offers a distinct advantage for a physically well-motivated and consistently defined measure of the local gravitational deformation, applicable even in regions far from filament spines. Our analysis uses the final MaNGA data release (2022ApJS..259...35A), which contains nearly 10,000 galaxies, nearly doubling the sample size of previous IFS-based studies and enabling a statistically robust detection of alignment signals. These advances allow us to probe the intrinsic connection between galaxy spin and the underlying cosmic tidal field, offering a new window into the physics of galaxy formation within the cosmic web.
@@ 2.1 MaNGA Galaxies
We utilize the IFS data from the complete release of the MaNGA survey, which is part of SDSS Data Release 17 (2015ApJ...798....7B; 2022ApJS..259...35A). All MaNGA data can be accessed via a Python-based tool called Marvin (2019AJ....158...74C). MaNGA employs a set of integral-field units, each consisting of between 19 and 127 fibers, to obtain spatially resolved spectra of nearby galaxies with a spectral resolution of $R\sim 2000$ and a wavelength coverage of 3600–10400 Å. The raw spectra are flux-calibrated and combined into 3D data cubes through the MaNGA Data Reduction Pipeline (DRP; 2016AJ....152...83L; 2021AJ....161...52L), and these data cubes are then processed by the MaNGA Data Analysis Pipeline (DAP; 2019AJ....158..160B; 2019AJ....158..231W) to derive various astrophysical quantities, including stellar and gas kinematics, emission-line fluxes, and spectral indices. The DAP performs full-spectrum fitting on the binned spectra, providing the line-of-sight velocity fields for both the stellar and ionized gas components.
@@
Our sample includes all galaxies observed in MaNGA, except those from four ancillary programs targeting nongalactic objects (DEEP_COMA, IC342, M31, and GLOBULAR_CLUSTER). We also exclude data flagged by the DAPQUAL bitmask as critical failures in either the DRP or DAP. Our parent sample consists of 10,058 unique galaxies, after removing duplicate entries in MaNGA. In this study, we include only disk galaxies (S0 and spiral types) classified in the MaNGA Visual Morphology Catalogue (2022MNRAS.512.2222V), resulting in a sample of 8199 galaxies.
@@
We measure the position angles (PAs) of the kinematic major axis, $\phi_{\mathrm{kin}}$, of MaNGA galaxies using the PaFit package (2006MNRAS.366..787K), which determines the axis that maximizes the match between the velocity field and its bi-antisymmetry model. The central positions and redshifts of the galaxies are adopted from the MaNGA catalog, which for most galaxies is identical to the NASA–Sloan Atlas (NSA; 2011AJ....142...31B). The MaNGA target selection strategy assigns 63 % of the main sample to reach 1.5 effective radii ($R_{e}$) and 37 % to reach $2.5\,R_{e}$ (2015ApJ...798....7B). For our analysis, we consistently use an elliptical aperture with a semi-major axis of $1.5\,R_{e}$ for each galaxy, with the aperture parameters defined by the NSA elliptical Petrosian photometry. We also exclude all bad pixels with inverse variance set to zero from the measurement. We assess the effect of the PA measurement uncertainties provided by PaFit through a Monte Carlo resampling test and find that their influence on the estimated spin alignment signal is negligible.
@@
The spin axis is defined as perpendicular to the kinematic major axis. The vector direction is determined by identifying the approaching and receding sides from the inverse-variance-weighted averages of line-of-sight velocities on either side of the spin axis. Following the method of 2007ApJ...671.1248L and 2011ApJ...732...99L, we compute the 3D unit spin vector, $\hat{\boldsymbol{j}}$, in spherical coordinates as
@@
where $\vartheta$ denotes the PA of the projected spin vector, measured eastward from north, ranging from 0 to $2\pi$. The inclination angle $i$ is estimated under a circular disk approximation from the observed axial ratio $(b/a)$ and the intrinsic axial ratio $p$, and is given by
@@
where we simply adopt $p=0.158$ to be consistent with 2021MNRAS.504.4626K.
@@
with $\alpha$ and $\delta$ representing the R.A. and decl. of each galaxy, respectively. Due to the sign ambiguity of the radial component $\hat{j}_{r}$, the 3D spin vector cannot be uniquely determined without additional assumptions, resulting in a twofold degeneracy (2011ApJ...732...99L; 2021NatAs...5..283M). While many previous studies arbitrarily selected one of the two vectors (e.g., 2021MNRAS.504.4626K; 2022MNRAS.516.3569B), we account for both in our analysis, which differ only by the sign of $\hat{j}_{r}$. We note that our results are not sensitive to either the sign of $\hat{j}_{r}$ or the value of $p$.
@@ 2.2 Reconstructed Tidal Field
We obtain the tidal field around the MaNGA galaxies from the Galaxy Environment for MaNGA Value Added Catalog (GEMA-VAC; 2015A&A...578A.110A; 2022ApJS..259...35A). The GEMA-VAC provides a characterization of large-scale structures based on the Exploring the Local Universe with the reConstructed Initial Density field (ELUCID) project (2016ApJ...831..164W). The present-day density field within the SDSS volume is reconstructed from the distribution of galaxy groups using the halo-domain method (2009MNRAS.394..398W; 2013ApJ...772...63W). The survey volume is partitioned into halo domains, and the mass distribution in each domain is modeled with $N$-body simulations. Redshift-space distortions are iteratively corrected using the line-of-sight component of the reconstructed linear velocity field.
@@
The density field is smoothed with a Gaussian kernel with a comoving scale of 2 $h^{-1}$Mpc. The three eigenvalues of $T_{ij}$, $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$, are obtained by diagonalizing $T_{ij}$ at the position of each galaxy. The corresponding eigenvectors are denoted by the major ($\hat{\boldsymbol{e}}_{1}$), intermediate ($\hat{\boldsymbol{e}}_{2}$), and minor ($\hat{\boldsymbol{e}}_{3}$) principal axes of $T_{ij}$, respectively. The major principal axis, $\hat{\boldsymbol{e}}_{1}$, corresponds to the direction of maximum matter compression. The eigenvalues can be used to classify the large-scale environment based on the number of positive eigenvalues: three positive eigenvalues indicate a cluster environment, two indicate a filament, one indicates a sheet, and none indicates a void. In our sample, a total of 6376 disk galaxies have corresponding tidal field data available for analysis. After excluding the small number lacking stellar mass or reliable PA, 6325 disk galaxies constitute the final analysis sample used throughout this work.
@@ 2.3 Alignment Angles
The spin–tidal alignment is quantified as the cosine of the angle, $|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|$ ($k=1,2,3$), where $\hat{\boldsymbol{j}}$ and $\hat{\boldsymbol{e}}_{k}$ represent the unit spin vector of each galaxy and the three principal axes of the local tidal tensor, respectively. In three dimensions, a random orientation yields an average of $|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|=0.5$, while parallel and perpendicular orientations correspond to $|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|=1$ and 0, respectively. Since the line-of-sight component of spin vectors cannot be uniquely determined in this study, each galaxy has two possible 3D spin orientations. In the edge-on view ($i=\pi/2$), the two vectors appear identical, while in the face-on view ($i=0$), they appear to point in opposite directions. To account for this ambiguity, we take the average of $|\hat{\boldsymbol{j}}\cdot\hat{\boldsymbol{e}}_{k}|$ from the two possible spin vectors to obtain a single alignment metric for each galaxy. This approach retains as much orientation information as possible while ensuring that the two spin vectors from the same galaxy are not treated as independent measurements. We also confirm that fixing the sign of $\hat{j}_{r}$ to either ++ or - leads to qualitatively identical alignment trends (see Appendix B), which is consistent with the findings of previous studies (e.g., 2021MNRAS.504.4626K; 2025ApJ...987L..30W).
The detection significance of the observed alignment is estimated from the 1000 random permutation samples in which the sky positions ($\alpha$ and $\delta$) of galaxies are randomly shuffled within the sample. For each random permutation, the 3D spin vectors are recalculated using the new sky coordinates through Equations (1)–(7). This approach provides the null hypothesis of random mutual orientation between galaxy spins and tidal fields, although it inevitably disrupts possible selection effects tied to galaxy sky positions and may therefore under- or overestimate the significance. We find, however, that alternative schemes (e.g., only randomizing $\phi_{\mathrm{kin}}$) do not alter the main conclusions of this work.